OpenStudy (sleepyjess):

Please let me know if I have done everything correctly, especially 2nd paragraph and then I need help and explanation on the last paragraph. TIA

OpenStudy (sleepyjess):

Data Months (x) Population P(x) 0 5 1 10 2 20 3 40 This exponential function measures growth monthly. Using complete sentences, how would I explain how to find the rate of growth every week. Originally, I thought I would calculate the Months (x) out to 12, which would be 20,480 and then divide that by 52, but the more I think about it, the more I am unconvinced that I am right. Next, I have to calculate when the population will reach 500, which would be [\frac{ 500 }{ 5 }=\frac{ 5(2^{x}) }{ 5 }\]\[100=2^{x}\]log(100) = (log)2^2\ log(100) = x (log)2 log(100)/log(2) = x x = 6.64385619 Last, the system I have to put the function into only works in logarithms of base 10. Using complete sentences, how would I explain converting my exponential function P(x) in a logarithmic one and then into a base 10 logarithm. This one I need help on.

OpenStudy (sleepyjess):

\[\frac{ 500 }{ 5 }=\frac{ 5(2^{x}) }{ 5 }\]

OpenStudy (sleepyjess):

Sorry... that what the equation should have been

OpenStudy (ranga):

Looking at the data we notice the population doubles every month. So the exponential function is 2^x. P(x) = A * 2^x when x = 0, P = 5 5 = A * 2^0 = A * 1. So A = 5 P(x) = 5 * 2^x where P is the population in month x.

OpenStudy (ranga):

If we want the Population as a function of w (where w represents week) then: P(w) = 5 * 2^(w/4) assuming 4 weeks in a month.

OpenStudy (ranga):

P(x) = 5 * 2^x When will the population reach 500? 500 = 5 * 2^x 2^x = 100 Take logarithm to the base 10 on both sides: x * log(2) = log(100) = 2 x = 2 / log(2) = 6.64 months

OpenStudy (sleepyjess):

I started there, but because there really isn't 4 weeks in a month, I went to the 52 weeks

OpenStudy (sleepyjess):

So I was correct on that part

OpenStudy (sleepyjess):

The 6.64 months

OpenStudy (ranga):

Yes. The population will reach 500 in 6.64 months.

OpenStudy (sleepyjess):

Okay, but what about the rate of growth every week. Wouldn't you have to calculate the growth at month 12 and then divide by 52 weeks?

OpenStudy (ranga):

This month to week conversion is an approximate one because months have: 28-31 days and the growth in month February will not be the same as the growth in March. But going the year route may be better than the 4 weeks to a month route.

OpenStudy (sleepyjess):

So would I be correct in my answer?

OpenStudy (ranga):

let me see. 52 weeks = 12 months 1 week = 12 / 52 months = 3/13 month P(x) = 5 * 2^x (where x is in months) OR P9w) = 5 * 2^(3w/13) (where w is in weeks)

OpenStudy (ranga):

P(w) = 5 * 2^(3w/13) (where w is in weeks)

OpenStudy (ranga):

Instead of dividing w by 4 to convert it to a month we are dividing w by 4.33333 to convert it to an average month.

OpenStudy (ranga):

1 week is 3/13th of a month w weeks will be (3w/13)th of a month P(x) = 5 * 2^x OR P(w) = 5 * 2^(3w/13)

OpenStudy (ranga):

For the last part, P = 5 * 2^x To convert an exponential function to a logarithmic function, we take log (to the base 10) on both sides of the equation: log(P) = log(5*2^x) = log(5) + log(2^x) = log(5) + x * log(2) log(P) = log(2) * x + log(5) If you want you can substitute decimal values of log(2) and log(5) or you can just leave it the way it is.

OpenStudy (sleepyjess):

Ok, I understand that but could it be done the way I was thinking?

OpenStudy (sleepyjess):

Thank you for the last part.

OpenStudy (ranga):

For the middle part, they actually want you to convert the function so that instead of population being a function of x (the month) they want the population to be a function of w (the week). So if I want to know what the population is in week 5, I should be able to put w = 5 in the formula and calculate P. The way you are doing, you are calculating the population after 1 year then dividing that by 52. That will be the averaging a years population growth equally among 52 weeks. That would assume a constant rate of growth week after week whereas this is an exponential growth.

OpenStudy (ranga):

If the question had been: Find the AVERAGE rate of growth per week (averaged over a one year period) then your method is correct.

OpenStudy (ranga):

Just noticed they want the RATE of growth, not just the population growth as a function of week! The general exponential growth function is: y = a * (1 + r)^x, where a is the initial value and r is the growth rate. P(w) = 5 * 2^(3w/13) P(w) = 5 * (2^(3/13))^w P(w) = 5 * 1.17346^w P(w) = 5 * (1 + 0.17346)^w 0.17346 or 17.346% is the growth rate.

OpenStudy (sleepyjess):

@ranga ... thank you so much... again... just got back to my computer and saw the rest of your post.

OpenStudy (ranga):

You are welcome.